OpenStudy (anonymous):

Help?! y=x+4 y^2+x^2=40,000

OpenStudy (anonymous):

and the question is where to they intersect? right?

OpenStudy (anonymous):

you can replace \(y\) by \(x+4\) and solve \[(x+4)^2+x^2=40,000\]

OpenStudy (anonymous):

\[x^2+8x+16+x^2=40,000\\ 2x^2+8x=39984\\ x^2+4x=19992\] etc

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw you can solve using the quadratic formula but it is probably easiest to complete the square in this case